OpenStudy (anonymous):

A boat traveled 294 miles downstream and back. The trip downstream took 14 hours. The trip back took 294 hours. Find the speed of the boat in still water and the speed of the current.

OpenStudy (anonymous):

boat speed = v1 current speed = v2 downstream the total speed is (v1+v2) on the trip back the total speed is (v1-v2) s = d/t d = 294 t = 14 while downstream t = 294 while "upstream" so now all you need is to solve this system of equations: v1+v2 = 294/14 v1-v2 = 294/294

OpenStudy (anonymous):

Can you show me in numbers how to solve the system of equations

OpenStudy (anonymous):

v1+v2 = 294/14 v1-v2 = 294/294 v1+v2 = 21 v1-v2 = 1 v1+v2 = 21 v1 = 1+ v2 (1+ v2) +v2 = 21 v1 = 1+ v2 2*v2 = 21 -1 v1 = 1+ v2 v2 = 20/2 = 10 v1 = 1 + v2 = 1+10 = 11 V2 = 10 mph V1 = 11 mph